A regression model relating x, number of salespersons at a branch office, to y, annual sales at the office (in thousands of dollars) provided the following computer output from a regression analysis of the data. Where n total=26.

a. Write the estimated regression equation (to whole number).

y=_____+_____x

b. Compute the F statistic and test the significance of the relationship at a .05 level of significance. (to 2 decimals)

F-value ____

p-value is _______, we _________ h0

c. Compute the statistic and test the significance of the relationship at a .05 level of significance. (to 2 decimals)

T - Stat =____

d. Predict the annual sales at the Memphis branch office. This branch employs 14 salespersons. (to whole number)

ANOVA
df SS MS F Significance F
Regression 6545.1
Residual
Total 8597.7
Coefficients Standard Error Stat P-value
Intercept 76 10.453
Number of Salespersons 43 5.799
Answer #1

a) y=76+43*x

b) F-value =MSR/SSE =(6545.1/1)/(2052.6/24)=76.53

p-value is =0.0000

we _reject___ h0

c)

t stat =43/5.799 =7.42

d)

Predicted annual sales =76+43*14=678